Draw a right rectangular prism that is
1) 2 units high, 3 units wide and 6 units deep

2) Explain how a net relates to a polyhedron and how you can calculate the surface of its corresponding polyhedron

Answer by RAY100(1637)

the prism is 2 x 3 x 6
.
net is 4 rectangles of 2 high, 3, 6, 3, 6 long
at rt end, two more rectangles come off of each 6 side, 3 high.
.
from net ,,,SA = 2 *(3 +6+3+6) +2(6*3) =72
,
from sides of prism, SA = 2(2*6) +2(2*3) + 2(3*6) = 72
.